OpenStudy (anonymous):

Which of these is a rule of integration 1/(a^(2) + x^(2))^(1/2) = arcsin(x/a), a>0 or 1/(a^(2) - x^(2))^(1/2) = arcsin(x/a), a>0 I'm leaning towards the second being a rule but I need clarification

OpenStudy (anonymous):

I dont need proof just clerificaiton but if you want to provide proof go nuts :)

OpenStudy (diyadiya):

\[\int\limits \frac{dx}{\sqrt{a^2-x^2}}= \sin^{-1} \frac{x}{a}+C\]

OpenStudy (anonymous):

nvm fine now

OpenStudy (anonymous):

thanks

OpenStudy (diyadiya):

Welcome:)